A hexagon is inscribed in a circle. If each side of the hexagon is 3, what is the area of the circle?

Answer by richwmiller(17219)

radius of hexagon=radius of circle=1/2*sqrt(3)*a
a=side
a=3
area circle=pi*r^2=pi*(1/2*sqrt3*3)^2=pi*1/4*3*9=27/4*pi=21.2